If (2a+1)^2+(4b-3)^2+(4c-5)^2=0 then find the value of a^3+b^3+c^3-3abc/a^2+b^2+c^2
With process

Grade:10

1 Answers

Arun
5 years ago
Dear student........ Given equation is satisfied only when a = -1/2.....b = ¾......C = 5/4....... Now......Put these values in the expression and you willbget your answer